THE HIGH-RESOLUTION INFRARED SPECTRUM OF DIBORANE

R.L. SAMS, T. BLAKE, AND S.W. SHARPE, Pacific Northwest National Laboratory, PO Box 999, Battelle Blvd. Richland, WA 99352; J.-M. FLAUD, Lab. Photophysique Moleculaire, CNRS, Universite Paris-Sud, Bat. 210, 91405 ORSAY, Cedex, France; \underline{W.J. LAFFERTY}, Optical Physics Division, NIST, Gaithersburg, MD 20899 and Instituto de Estructura de la Materia, C.S.I.C., Serrano 121, 28006 Madrid, Spain..

We are in the process of recording and analyzing all of the infrared active bands of B_{2}H_{6}, the simplest of the hydrogenbridged borane compounds, with the goal of determining the frequencies of the IR forbidden modes in this very symmetric (D_{2h}) molecule by means of anharmonic interactions. The spectra were obtained on a Bruker 120HR FT spectrometer with a spectral resolution of 0.002 cm^{−1}. The sample contained boron in natural abundance (∼80% ^{11}B; 20% ^{10}B) and, therefore, the abundance of ^{11}B_{2}H_{6}, ^{10}B^{11}BH_{6} and ^{11}B_{2}H_{6} was 64%, 32% and 4% respectively. With the resolution available, the spectra of the two most abundant isotopomers were readily assignable. At this point, the ν_{14} (c-type) and ν_{17} (a-type) bands have been assigned for both the ^{11}B_{2}H_{6} and ^{11}B^{10}BH_{6} isotopic species and ground state and upper state constants obtained. Both bands exhibit small perturbations. A mixed r_{s}-r_{0} structure for diborane has been obtained.